OpenStudy (anonymous):

Rewrite using scientific notation: 345,000

OpenStudy (tommo_lcfc):

3.45 * 10^5

OpenStudy (anonymous):

3.45X10^5

OpenStudy (anonymous):

all wrong it's 3.45x10^5

OpenStudy (tommo_lcfc):

The answers are all the same, it does not matter about whether we put *, x or X. It means the same thing. I know you're joking :)

Parth (parthkohli):

= 3.45 * 100000 = 3.45 * 10^5